How long will it take an athlete to run 300m at a gradient of 1:10 if he can run 300m in 60 sec on a flat surface. Weather is a constant.
1
Expert's answer
2019-10-01T12:45:26-0400
"s=0.5at^2"

"a=\frac{2s}{t^2}"

"a'=a-g\sin{\alpha}=a-0.1g"

"t'=\sqrt{\frac{2s}{a'}}=\sqrt{\frac{2s}{a-0.1g}}=\sqrt{\frac{2s}{\frac{2s}{t^2}-0.1g}}"

"t'=t\sqrt{\frac{1}{1-\frac{0.1gt^2}{2s}}}=60\sqrt{\frac{1}{1-\frac{0.1(9.8)(60)^2}{2(300)}}}=93\ s"